### v2

# BarcodeBERT: Transformers for Biodiversity Analyses

Abstract
In the global challenge of understanding and characterizing biodiversity, short species-specific genomic sequences known as DNA barcodes play a critical role, enabling fine-grained comparisons among organisms within the same kingdom of life. Although machine learning algorithms specifically designed for the analysis of DNA barcodes are becoming more popular, most existing methodologies rely on generic supervised training algorithms. We introduce BarcodeBERT, a family of models tailored to biodiversity analysis and trained exclusively on data from a reference library of 1.5 M invertebrate DNA barcodes. We compared the performance of BarcodeBERT on taxonomic identification tasks against a spectrum of machine learning approaches including supervised training of classical neural architectures and fine-tuning of general DNA foundation models. Our self-supervised pretraining strategies on domain-specific data outperform fine-tuned foundation models, especially in identification tasks involving lower taxa such as genera and species. We also compared BarcodeBERT with BLAST, one of the most widely used bioinformatics tools for sequence searching, and found that our method matched BLAST’s performance in species-level classification while being 55 times faster. Our analysis of masking and tokenization strategies also provides practical guidance for building customized DNA language models, emphasizing the importance of aligning model training strategies with dataset characteristics and domain knowledge. The code repository is available at https://github.com/bioscan-ml/BarcodeBERT.
1 Introduction
The task of estimating and understanding biodiversity on our planet remains a monumental challenge, as traditional methods of taxonomic analysis often struggle to keep pace with the rate of discovery and identification of new species. In this context, the search for highly expressive, short standardized genomic regions containing meaningful taxonomic information (DNA barcodes) has become prominent in biodiversity research over the past two decades (Hebert et al., 2003; Miller, 2007; Hebert et al., 2016; Srivathsan et al., 2021). Specifically, a 658-base-pair-long fragment of the Cytochrome c Oxidase Subunit I (COI) gene (Lunt et al., 1996) has emerged as the de facto DNA barcode for kingdom Animalia (Dopheide et al., 2019) and has proven effective in addressing inherent taxonomic challenges. Particularly, barcodes can be used for fast and accurate queries to categorize novel specimens into existing taxa. Furthermore, in the absence of clear species boundaries, they can be used to systematically separate specimens into groups of closely related organisms. These clusters, known as operational taxonomic units (OTUs) correspond to groups of similar specimens and can be labelled using e.g. a Barcode Index Number (BIN) (Ratnasingham & Hebert, 2007). As it is defined systematically, such a BIN system overcomes ambiguities in traditional species labelling and thus accelerates biodiversity research. Among the numerous taxonomic groups to which DNA barcoding is applicable, invertebrates, particularly arthropods, stand out as an incredibly diverse and taxonomically complex group (Basset et al., 2012), making them the focus of many methodological studies (Badirli et al., 2021, 2023; Gong et al., 2024). The diversity and taxonomic richness of this group require specialized algorithmic approaches that can capture the taxonomic structure of the data. Consequently, biodiversity researchers are increasingly turning to machine learning methods, including convolutional neural networks (CNNs) (Badirli et al., 2021) and transformer models (Gong et al., 2024), to scale taxonomic classification of arthropods and accelerate species discovery.
Transformer-based models, pretrained at scale with self-supervised learning (SSL), also referred to as “foundation models,” have found applications across diverse domains thanks to their effectiveness in learning from large unlabelled datasets (Brown et al., 2020; Touvron et al., 2021). These models are often task-agnostic and can perform well on a variety of downstream tasks after fine-tuning. Despite their success in other domains, their application for taxonomic identification using DNA barcodes has not yet been extensively explored. Moreover, most DNA-based foundation models primarily target human chromosomal DNA sequences (Zhou et al., 2023; Dalla-Torre et al., 2023; Ji et al., 2021), making them suboptimal for barcode data due to domain shift between the two data types. In particular, DNA barcodes for animals stem from a specific region in mitochondrial DNA, and patterns learned from other genomic regions may be irrelevant for taxonomic classification.
We here aim to unlock the potential of transformer-based architectures for taxonomic identification of arthropod barcodes, providing insights that extend beyond broad, foundation-style approaches. We address the previously mentioned issues (i.e. the taxonomic complexity of arthropods, and the lack of specialized transformer models trained on DNA barcodes) by adopting a semi-supervised learning approach, followed by fine-tuning on high-quality labelled barcode data, demonstrating the value of targeted model development for specialized applications. We propose BarcodeBERT, a self-supervised method that leverages a reference library of 1.5M invertebrate barcodes (deWaard et al., 2019) and a masked language model (MLM) training strategy to effectively compute meaningful embeddings of the data, facilitating successful species-level classification of insect DNA barcodes in general scenarios. In addition to the classification of known species, our pretrained models can be used to generate embeddings for sequences from unseen taxa, enabling non-parametric classification at higher levels of the taxonomic hierarchy.
To summarize our contributions, we first investigate the impact of pretraining using a large and diverse DNA barcode dataset (1 million sequences, from more than 17,000 species, across 6,700 genera) on generalization to other downstream tasks. Second, we compare BarcodeBERT against several baselines such as pretrained DNA foundation models (DNABERT (Ji et al., 2021), DNABERT-2 (Zhou et al., 2023), DNABERT-S (Zhou et al., 2024), the Nucleotide Transformer NT (Dalla-Torre et al., 2023), and HyenaDNA (Nguyen et al., 2023)), a CNN baseline following the architecture introduced by Badirli et al. (2021), and the widely used alignment-based method BLAST (Altschul et al., 1990). Third, our study provides actionable insights regarding tokenization strategies, optimal masking ratios, and the importance of application-specific pretraining for DNA language models.
Overall, BarcodeBERT outperforms all other foundation models in supervised species classification, matching BLAST’s accuracy while being 55 times faster and more scalable. Moreover, a linear classifier trained on BarcodeBERT embeddings has $\sim$6higher species classification accuracy than the top-performing foundation model in this task. Lastly, the same embeddings can also be used for accurate genus classification using similarity searches, outperforming the top-performing foundation model by $\sim$30%.
2 Related Work
The exponential growth of genomic datasets with the advent of high-throughput sequencing has both demanded and enabled a surge in classification tools for DNA sequences. Such tools are essential for large-scale biodiversity studies, where algorithmic approaches can expedite the taxonomic categorization of novel specimens. One intuitive approach is to embed sequences into a vector space where geometric distances approximate taxonomic similarities (Corso et al., 2021). This allows for rapid comparisons between newly sequenced and labelled DNA, enabling accurate taxonomic assignments.
Many machine learning approaches, particularly in the area of representation learning, have demonstrated considerable potential in biodiversity analyses as they can embed raw DNA data into an expressive lower dimensional space. Transformer-based models (Vaswani et al., 2017), known for their ability to capture complex patterns within sequential data, have shown exceptional performance in various representation learning tasks across domains, either with or without supervision (Brown et al., 2020; Touvron et al., 2021; Caron et al., 2021). These models are especially effective in learning from vast unlabelled datasets, making them ideal candidates for the analysis of genomic data, where obtaining high-quality annotations remains challenging.
There has been a growing number of self-supervised learning-based DNA language models proposed recently, most of which are based on the transformer architecture and trained using the masked language model (MLM) objective. The first foundational model in this space, DNABERT, utilizes a BERT-based transformer architecture along with $k$-mer tokenization for genome sequence prediction tasks. Following DNABERT, other models have emerged, including the Nucleotide Transformer (Dalla-Torre et al., 2023), GENA-LM (Fishman et al., 2023), and HyenaDNA (Nguyen et al., 2023). While each model varies in architectural details, tokenization methods, and training data, their reliance on SSL and the MLM objective for pretraining remains a constant. HyenaDNA is a unique entry in this space as it uses a state-space model (SSM) based on the Hyena architecture (Poli et al., 2023) and trains it for next-token prediction (a causal MLM).
The landscape of machine learning models specifically tailored for DNA barcodes is less developed. A recent study (Badirli et al., 2023) proposes a Bayesian framework based on CNNs which, when combined with visual information, achieves high accuracies in species-level identification of seen species and genus-level inference of novel species in a dataset of $\sim$32,000 insect DNA barcodes. This method uses supervised learning to compute meaningful embeddings that can be used as side information in a two-layer Bayesian zero-shot learning framework. Transformer methods have been introduced for the classification of fungal Internal Transcribed Spacer sequences without any self-supervision (Romeijn et al., 2024).
Although there has been a growing number of SSL-based DNA language models proposed in the recent literature, our findings indicate that models pretrained on a diverse set of non-barcode DNA sequences underperform on downstream barcode tasks. This suggests that general DNA foundation models may struggle with the domain-specific characteristics of barcode data. In this study, we leverage barcode-specific training to improve both species-level classification accuracy and generalization to other taxonomic ranks. By grounding our approach in targeted data and architectural choices, we seek to advance the utility of machine learning in biodiversity research, moving beyond general off-the-shelf models trained to classify specimens into known taxa. Distinctly, our specialized models are not only capable of classifying known species but also can be used for taxonomic classification for species that are not present in the training set.
3 Methods
In this section, we outline the key elements of our methodology. We begin with a detailed account of our datasets and data processing pipeline. We then describe the architectures and hyperparameters used in the development of BarcodeBERT.
3.1 Dataset
We use a reference library for Canadian invertebrates (deWaard et al., 2019) for model training and testing, comprised of 1.5 M DNA samples from the Barcode of Life Data system (BOLD) (Ratnasingham & Hebert, 2007). The dataset was further pre-processed and subdivided as described below.
3.1.1 Data pre-processing
To ensure data integrity and consistency, we performed a series of pre-processing steps over this dataset. First, empty entries were removed. Then, following standard practices (Dalla-Torre et al., 2023), IUPAC ambiguity codes (non-ACGT symbols), including alignment gaps, were uniformly replaced with the symbol N. Duplicate sequences were removed to avoid redundancy and increase the complexity of the training and pretraining tasks. Sequences with trailing N’s were truncated. Finally, sequences falling below 200 base pairs or exhibiting over 50% N characters were excluded.
3.1.2 Data partitioning
After pre-processing, 965,289 barcode sequences from 17,464 invertebrate species, across 6,712 genera were obtained. The dataset was divided into three distinct partitions for different training and evaluation purposes: (i) Seen: This partition is intended for supervised learning pipelines, particularly to evaluate the model’s ability to classify specimens from well-represented taxa. Comprised only of samples labelled to species-level, it includes 67,267 barcodes from 1,653 arthropod species, across 500 different genera, with each species represented by at least 20 and at most 50 barcodes. The partition is further split into training (70%), testing (20%), and validation (10%) subsets. (ii) Unseen: This test partition was sampled to evaluate the models in real-world conditions where specimens from underrepresented species are frequently obtained. It only contains barcodes from “rare” species with fewer than 20 barcodes in the full reference dataset. Specifically, this partition contains 4,278 barcodes from 1,826 arthropod species, none of which are present in any other partition. Moreover, this partition contains all 500 genera labels present in the Seen partition, with up to 20 barcodes sampled per genus. The label distribution shifts are shown in Figure 1, with the Seen partition reflecting the overall dataset’s distribution and the Unseen partition exhibiting a greater diversity of rare genera. (iii) Pretrain: This partition contains the remaining 893,744 barcode sequences from 14,794 invertebrate species across 6,679 genera. Note that only 35% of the sequences in this partition contain full taxonomic annotations up to the species level. See Appendix A for more details on dataset composition.
Figure 1: Distribution of orders in the Seen (left) and Unseen (right) partitions of the dataset. Icons: CC BY-SA, Wikimedia; Pro Content license, Canva.
3.2 Proposed method: BarcodeBERT
Inspired by Bidirectional Encoder Representations from Transformers (BERT)-like models, which convert sequence inputs into meaningful embedding vectors, BarcodeBERT is designed to encode DNA barcodes into informative embedding vectors for fast and effective comparisons. This architecture’s main building block is the transformer layer, with multi-head attention units playing a crucial role in capturing positional dependencies within each input sequence. Our model features four transformer layers, each with four attention heads, enabling a robust representation of the DNA barcode data while maintaining a manageable number of hyperparameters. Figure 2 shows the details of BarcodeBERT architecture.
Figure 2: Architecture of BarcodeBERT, a transformer-based model employing a self-supervised learning strategy. The model is trained on non-overlapping $k$-mers from DNA barcode sequences. Any $k$-mer containing a character that is not in the nucleotide vocabulary is replaced by the [UNK] token. Pretraining involves masking out certain input parts using the [MASK] token and predicting these masked elements using a linear classification head. During training, the model selects a random offset ($0\leq$ offset $<k$) for each sequence and begins tokenization from that position. This helps create more robust embeddings and increases resilience to potential mutations.
Before being fed as input to the model, each barcode is split into a sequence of tokens. After evaluating two of the most common tokenization strategies for DNA sequences, Byte Pair Encoding (BPE) (Zhou et al., 2023; Romeijn et al., 2024) and $k$-mer tokenization (Ji et al., 2021; Dalla-Torre et al., 2023), we selected non-overlapping $k$-mer tokenization for BarcodeBERT (see the Ablation Studies section for more details). The token vocabulary includes all possible $k$-mer combinations derived from the nucleotide alphabet {A,C,G,T}, supplemented by two special tokens: [MASK] and [UNK]. The [MASK] token is utilized for masking $k$-mers during the pretraining phase, and $k$-mers containing any symbol that is not present in the nucleotide alphabet are assigned the [UNK] token. This results in a vocabulary size of $24^{k}+$ 2.
A limitation of this tokenization strategy is its sensitivity to frame shifts. For example, the $k$-mer representation of the sequence GATCGA differs entirely from that of CGATCGA, even though the sequences differ by only a one-nucleotide shift. To address this issue and make our model robust to frame shifts that may occur in practice, we introduce a data augmentation step by randomly offsetting the sequence by a value ($0\leq$ offset $<k$) during pretraining to improve generalization. Before tokenization, DNA barcodes are either padded or truncated at 660 nucleotides to ensure coverage of the barcode region in the COI gene. Finally, the tokenized sequences are fed to the model and encoded into a sequence of $768$-dimensional vectors.
Following self-supervised training, our model produces a whole barcode-level embedding vector by applying global average pooling over the sequence of $d$-dimensional output vectors, ignoring padding and any special tokens. During inference, the pipeline mirrors the training setup without the random offset: DNA barcodes are tokenized into non-overlapping $k$-mers and passed through the model, generating embeddings that capture meaningful taxonomic information across the entire sequence. BarcodeBERT is implemented using PyTorch and the Hugging Face Transformers library. During training, we focused exclusively on masked token prediction, masking 50% of the input tokens and optimizing the network with a cross-entropy loss. We optimize the model parameters by gradient descent using the AdamW (Loshchilov & Hutter, 2017) optimizer with weight decay set to $1\times 10^{-5}$ and a OneCycle schedule with maximum learning rate of $1\times 10^{-4}$. Additionally, we performed experiments across different $k$-mer lengths ($2\leq k\leq 8$) to observe the impact of $k$-mer length on embedding quality.
4 Experiments
In this section, we present our evaluation framework and evaluate the performance of BarcodeBERT against the baseline models across several tasks. Additionally, we present a series of ablation studies to justify our design choices and analyze the impact of key hyperparameters on the model’s performance.
4.1 Experimental setup
To explore the applicability of our model for DNA barcode-based biodiversity analyses, we employ different SSL evaluation strategies (Balestriero et al., 2023) and contrast its performance against the baselines. First, we evaluate our models in a “closed-world” setting where the goal is to classify DNA sequences into known taxa.
Fine-tuning. Pretrained models are fine-tuned on the training subset of the Seen partition and evaluated on the test subset. This task assesses the ability of models to perform species-level classification with full access to labelled training data.
Linear probing. To evaluate the quality of pretrained embeddings, the backbone of the models is frozen, and a linear classifier is trained on the training subset of the Seen partition. The final classifier is evaluated on the test subset, providing insights into the effectiveness of the embeddings without extensive task-specific training.
1-NN probing. This task evaluates model generalization to new species within known genera. Using cosine similarity, we perform 1-NN probing at the genus level with the training subset of the Seen partition as the reference set and the Unseen partition as the query set.
Second, our models are evaluated in an “open-world” setting where the goal is to group sequences, including those from unknown species, based on shared features.
BIN reconstruction. We merge the test subset of the Seen partition with the Unseen partition and evaluate the model’s ability to reconstruct Barcode Index Numbers (BINs) using embeddings generated without fine-tuning in a zero-shot clustering (ZSC) task (Lowe et al., 2024). This task assesses how well the embeddings capture the hierarchical structure of taxonomic relationships, including rare or unclassified species.
Finally, we evaluate the utility of learned DNA embeddings as auxiliary information in multi-modal learning.
Bayesian zero-shot learning. We selected a species-level image classification task using the INSECT dataset (Badirli et al., 2021). This is a small multimodal dataset designed for zero-shot classification of images from unseen species using DNA as auxiliary information. DNA embeddings generated by the models are paired with pre-extracted image features to classify species in a zero-shot setup. We evaluate both embeddings from pretrained and fine-tuned models on the species classification task from the INSECT dataset. Following Badirli et al. (2021), the Bayesian zero-shot learning (BSZL) framework uses image features as priors and DNA embeddings as side information. For unseen species, the $K$-nearest seen species in the DNA embedding space are used to define local priors, allowing the Bayesian model to generate posterior predictive distributions for unseen categories. To ensure a fair comparison with prior work, image features are pre-extracted using ResNet-101 (He et al., 2016). Hyperparameter tuning for the Bayesian model is performed using the same grid search space as in Badirli et al. (2021).
4.2 Results
In this section, we describe our results on two evaluation tasks: DNA-specific tasks, designed to assess model performance in both open- and closed-world taxonomic settings; and zero-shot image classification using DNA embeddings.
4.2.1 DNA-specific categorization tasks
Our evaluation, presented in Table 1, compares several models across species-level and genus-level DNA-specific categorization tasks (fine-tuning, linear probing, 1-NN probing, BIN reconstruction through ZSC). For species-level classification, we performed a BLAST search for reference and obtained a 99.7% classification accuracy after the selection of the best hit. The performance of all fine-tuned deep learning-based models is comparable with this baseline, and all transformer models outperform the CNN model as well. DNABERT-2, DNABERT-S and BarcodeBERT achieved nearly identical accuracies over 99.7%. Notably, only BarcodeBERT continues to closely match BLAST’s performance using a linear classifier, highlighting its strength in encoding meaningful features from raw data. In genus-level 1-NN probing, BarcodeBERT achieves the highest accuracy (78.5%) among the deep learning-based models, demonstrating a superior ability to generalize across taxonomic levels. BLAST, however, performs best in this task with 83.9%. This result indicates that, without fine-tuning, BarcodeBERT captures coarser taxonomic distinctions but is limited in representing the full hierarchical taxonomic structure as illustrated in Figure 3. The ZSC task provides additional insights into the model’s understanding of the hierarchical taxonomic structure. High performance in ZSC alone indicates a learned representation’s ability to finely distinguish between closely related clusters (BINs) without necessarily capturing the higher-level taxonomy. In contrast, strong 1-NN performance at higher taxonomic levels but lower ZSC accuracy suggests that the model understands the overall topology of the hierarchical taxonomic structure, even if it lacks the granularity needed for precise clustering. DNABERT and BarcodeBERT exhibit this distinction, with BarcodeBERT achieving a more balanced performance across tasks, making it the more versatile model for comprehensive DNA barcode analysis.
Table 1: Classification accuracy of DNA barcode models under different SSL evaluation strategies and different efficiency metrics. The baselines are divided into three groups: alignment-based techniques, BLAST; a deep learning-based non-SSL CNN baseline; and off-the-shelf DNA foundation models pretrained on non-barcode data. These are compared against BarcodeBERT, which is specifically pretrained on DNA barcode-based datasets. For BarcodeBERT we used the best configuration of $k=4$, with 4 attention heads, and 4 layers (4-4-4). Some models supported variable stride length; for these, the numbers in parentheses are the optimal $k$-mer values that yield the best results. We also show the throughput-per-second (TPS) of the encoders, and the total duration of the classification tasks. Numbers in boldface indicate the best result across each task, and underlined indicates second place.
Species-level acc (%) of seen species Genus-level 1-NN probe of unseen species BIN reconstruction accuracy (%)
Model #Param. TPS (seq/s) Finetuned Linear probe Dur (s) Acc (%) Dur (s) ZSC probe
BLAST N/A N/A 99.7 1495 83.9 602 N/A
CNN encoder 1.8 M 934 98.2 51.8 13 47.0 55 26.8
DNABERT 88.1 M 50 ($k$=6) 99.5 ($k$=4) 47.1 248 ($k$=6) 48.1 1021 79.3
DNABERT-2 118.9 M 134 99.7 87.2 101 23.5 381 38.1
DNABERT-S 117.1 M 134 99.7 93.1 101 30.6 381 62.7
HyenaDNA-tiny 1.6 M 1167 99.2 93.5 11 37.5 44 25.8
Nucleotide Transformer 55.9 M 95 99.5 65.1 140 40.1 536 22.4
BarcodeBERT (4-4-4) 29.1 M 484 99.7 99.0 27 78.5 108 73.2
BLAST is a deterministic algorithm without any learning component (see Appendix D for details). Consequently, species classification accuracy does not correspond to fine-tuning or linear probing, and it is only included in the table for reference.
Two efficiency measurements are included: throughput, defined exclusively for deep-learning-based models as the number of sequences processed per second, and total runtime for classification pipelines to ensure a fair comparison with alignment-based baselines. In terms of throughput, HyenaDNA(tiny) showcases the capabilities of state space models, demonstrating high throughput with fewer parameters. However, its classification performance is lower compared to BarcodeBERT and DNABERT-2. In total run time, our results indicate that subquadratic methods like the CNN baseline and HyenaDNA perform genus-level similarity searches (1-NN probe) 13× faster than BLAST, while BarcodeBERT is 5.5× faster than BLAST and outperforms other transformer models at this task. For species-level classification pipelines that do not include the computation of the training embeddings, transformer-based models demonstrate clear advantages over traditional baselines in terms of running time. Notably, BarcodeBERT, with a moderate parameter count, matches BLAST’s high classification accuracy ($99.7\%$) with a 55× faster running time, thus providing a well-rounded option for large-scale DNA barcode applications. All efficiency experiments were conducted using an Intel(R) Xeon(R) CPU @ 2.20GHz processor and a Quadro RTX 6000 GPU.
Figure 3: Comparison of different DNA foundation models on the task of 1-NN probing at different taxonomic levels. The query set contains DNA barcodes from species not present in the key set, and none of the models have undergone fine-tuning.
4.2.2 Zero-shot image classification using DNA embeddings
We use the Bayesian zero-shot learning task to evaluate the quality of the DNA feature embeddings, by assessing their effectiveness when used as side information for classifying images to species on the INSECT dataset. We consider the embeddings directly from the pretrained models and also after fine-tuning. The accuracy for seen and unseen test species and the harmonic mean are presented in Table 2. Even without fine-tuning, BarcodeBERT substantially outperforms DNABERT and DNABERT-2 on unseen species, regardless of whether they had been fine-tuned previously. BarcodeBERT achieves similar performance to the reported baseline CNN results (Badirli et al., 2021) and improves on the harmonic mean score by 1.2% and unseen accuracy by 1.9%, respectively. Our results demonstrate that in the zero-shot learning task of predicting insect species, employing BERT-like models that have also been trained on insect DNA barcodes as DNA encoders can improve performance over CNNs and general DNA foundation models.
Table 2: Evaluation of DNA barcode models in a Bayesian zero-shot learning task on the INSECT dataset. The pretraining and fine-tuning data source is indicated by the respective DNA type, and ‘–’ signifies the absence of training for that type. We also indicate the most specific taxon subset. For the baseline CNN encoder, we report the original paper result (left) and reproduced result (right). Numbers in boldface indicate the best result across each task, and underlined second best.
Data sources Species-level acc (%)
Model SSL pretraining Fine-tuning Seen Unseen Harmonic Mean
CNN encoder Insect 38.3 / 39.4 20.8 / 18.9 27.0 / 25.5
DNABERT Human 35.0 10.3 16.0
DNABERT Human Insect 39.8 10.4 16.5
DNABERT-2 Multi-species 36.2 10.4 16.2
DNABERT-2 Multi-species Insect 30.8 8.6 13.4
BarcodeBERT (ours) Invertebrates 31.6 20.0 24.5
BarcodeBERT (ours) Invertebrates Insect 38.8 15.3 22.0
4.3 Ablation Studies
We study the impact of the different components involved during pretraining. We use the terminology context tokens for tokens that are left unchanged to provide context to the model during pretraining and the terminology substitution tokens for tokens that will be changed for the masked language modelling task. We consider different strategies to calculate the loss of each group separately. The loss associated with predicting contextual tokens is referred to as the “context component” of the loss [^1] Although predictions in the foundation model literature are typically restricted to substitution tokens, we extend this to include context token predictions, maintaining the terminology to explore their potential utility in input reconstruction similar to non-masked autoencoding objectives., while the loss related to predicting substitution tokens is the “substitution component”. By assigning different weights to these two loss components, we sought to observe how these adjustments would affect both training and evaluation. In particular, we define $w_{s}$ as the penalty weight given to the substitution component of the loss. Note that $1-w_{s}$ is always the weight of the context component of the loss.
4.3.1 Substitution token rate
To examine how varying the substitution token ratio ($r_{s}$) affects performance, we tested several ratios, keeping the model architecture (4 attention heads, 4 layers), tokenization ($k=4$) and substitution penalty weight ($w_{s}=1$) constant. Table 3 shows that species-level classification performance remains consistently high across substitution rates, peaking at 99.67% accuracy with 45% and 50% substitution tokens. Linear probe results align closely, reaching the highest accuracy of 99.02% at the 50% substitution rate. For genus-level 1-NN probing of unseen species, the 50% substitution rate yields the best accuracy at 78.47%, suggesting that this rate provides a balance that strengthens the model’s ability to generalize to new taxa. Lower substitution rates show slightly reduced generalization, while a 60% rate begins to degrade performance, indicating that 50% is the optimal value for $r_{s}$.
4.3.2 Weight of the substitution component of the loss
Building on the fact that predicting context tokens is inherently easier than predicting substitution tokens for LLMs, we investigated how adjusting the penalty weights between these two tasks affects the performance of the model. For this purpose, we experimented with different penalty weights assigned to the substitution component of the loss ($w_{s}$). Table 4 provides the accuracy for genus-level 1-NN probing of unseen species across different values for $w_{s}$ in combination with four $k$-mer sizes (2, 4, 6, and 8) and Byte-Pair Encoding (BPE) tokenizer obtained from DNABERT-2. Alternative BPE tokenizers that specifically fit our data are investigated later in this section. We kept the architecture (4 layers, 4 attention heads) and substitution token rate ($r_{s}=50\%$) constant. As shown in Table 4, the optimal performance across all $k$-mer sizes was achieved with a $w_{s}$ of 1.0, where the highest accuracy, 78.47%, was observed with $k=4$.
Table 3: Classification accuracy over the different substitution token ratios $r_{s}$, while keeping constant all of the model architecture (4-4), the value of $k=4$ during tokenization and the penalty weight for the substitution component of the loss ($w_{s}=1$). Numbers in boldface indicate the highest accuracies and the $\rightarrow$ highlighted $\leftarrow$ value shows the selected optimal parameter.
Substitution token ratio (%) Species-level acc (%) of seen species Genus-level acc (%) of unseen species
Finetuned Linear probe 1-NN probe
15 98.95 98.95 75.15
30 98.79 98.79 74.24
45 99.67 98.54 74.42
$\rightarrow$ 50 $\leftarrow$ 99.67 99.02 78.47
60 99.62 98.45 77.56
Table 4: Genus-level accuracy for 1-NN probing of unseen species with varying penalty weight assigned to the substitution component of the loss ($w_{s}$). The model architecture remains fixed (4-4), with substitution token ratio $r_{s}$ = 50%. Two tokenizers were tested: a $k$-mer tokenizer with $k$-mer sizes of 2, 4, 6, and 8, and a BPE tokenizer used in DNABERT-2 with a fixed vocabulary size of 4096. Note that the weight of the context component of the loss always equals $1-w_{s}$. The number in boldface indicates the overall best accuracy, underlined the best per tokenizer, and $\rightarrow$ highlighted $\leftarrow$ the selected optimal parameter.
Genus-level acc (%) of
unseen species with $1$-NN probe
Loss weight ($w_{s}$) $k\!=\!2$ $k\!=\!4$ $k\!=\!6$ $k\!=\!8$ BPE
0.2 64.18 76.06 75.15 71.15 70.57
0.5 66.47 74.98 76.62 71.22 70.34
0.8 68.84 76.71 74.66 73.33 69.40
0.9 69.51 77.16 76.06 72.23 67.48
$\rightarrow$ 1.0 $\leftarrow$ 76.92 78.47 75.74 75.62 69.85
Our experiments indicate that focusing the loss penalty on the harder task of predicting substitution tokens, while not penalizing the easier task of predicting context tokens, yields the best accuracy. This aligns with observations in other foundation models, such as BERT and DNABERT (see Appendix C for more details).
4.3.3 Tokenization strategies
Table 5: Genus-level accuracy for unseen species with different tokenizers, various model sizes, fixed weight for the substitution component of the loss function ($w_{s}\!=\!1$) and a substitution token ratio $r_{s}\!=\!50\%$. Two types of tokenizers were tested: a $k$-mer tokenizer with $k$-mer sizes of 2, 4, 6, and 8, and five different versions of BPE tokenizers. BPE tokenizer from DNABERT-2 has a fixed vocabulary size of 4096, while BPE tokenizers trained on our dataset have vocabulary sizes ($v$) of 4096, 1024, 256, and 128. Numbers in boldface indicate the best result across each architecture.
$k$-mer tokenizer DNABERT-2 BPE BarcodeBERT BPE
Model size $k\!=\!2$ $k\!=\!4$ $k\!=\!6$ $k\!=\!8$ $v\!=\!4096$ $v\!=\!4096$ $v\!=\!1024$ $v\!=\!256$ $v\!=\!128$
4 layers, 4 heads 76.92 78.47 75.74 75.62 69.85 66.88 68.58 66.57 63.42
6 layers, 6 heads 71.46 76.95 76.04 76.60 70.17 67.30 66.95 63.49 60.61
12 layers, 12 heads 74.71 70.17 70.80 75.81 68.68 67.79 62.39 56.94 54.09
We evaluated the BPE tokenizer, which generates variable-length tokens based on character co-occurrence frequencies (Sennrich et al., 2016). BPE was designed for subword tokenization to overcome fixed vocabulary (Sennrich et al., 2016) by compressing sequences to reduce the information required for representation (Gallé, 2019). Unlike overlapping $k$-mers, BPE has an advantage for masked DNA language models (Zhou et al., 2023), since it avoids information leaks from adjacent masked tokens. We used DNABERT-2’s BPE tokenizer, trained on 2.75 billion nucleotide bases from the human nuclear genome and 32.49 billion bases from 135 species across various kingdoms. Also, we trained custom BPE tokenizers with varying vocabulary sizes using our DNA barcode dataset and evaluated across various training scenarios on the genus-level 1-NN probing task. Based on results in Table 4, we selected a loss weight of 1.0 since it yielded the best results. Similarly, our setup followed from Table 4, as we maintain a 50% substitution token ratio ($r_{s}$). All the results for the various configurations can be found in Table 5.
Model size. We report genus-level 1-NN probing accuracy on the unseen data for DNABERT-2 BPE and BarcodeBERT BPE tokenizers across three model sizes. Models using DNABERT-2’s BPE showed no significant accuracy changes across different sizes. However, for BarcodeBERT BPE, we see that increasing the model size reduces the accuracy across different vocabulary sizes, possibly due to overfitting.
Vocabulary size. We trained new BPE tokenizers with varying vocabulary sizes ($v$) to evaluate their impact on BarcodeBERT BPE performance. Unlike $k$-mer tokenizers, BPE sequence lengths vary with input composition and require padding ([UNK]) or truncation to a maximum length before tokenization. Since smaller vocabularies ($v$) produce longer sequences (see Table 11), we set maximum lengths of 128 for $v=4096,1024$ and 256 for $v=256,128$. Reducing the vocabulary size from 4,096 to 128 consistently decreases accuracy across all model sizes.
Comparing $k$-mer with BPE. We find that $k$-mer tokenizers outperform BPE tokenizers in all model configurations. We hypothesize that this is likely due to three reasons. First, DNA barcode sequences are too short to benefit from the BPE compression. Second, BPE is sensitive to minor variations such as single-character substitutions, which is unsuitable for DNA datasets with single-nucleotide mutations in arbitrary positions. In other words, BPE tokenization varies greatly for similar sequences with small Hamming distances, while $k$-mer tokenization remains consistent, as a single-nucleotide substitution affects only one token (see Table 11). Third, although BPE handles small sequence alignment shifts better than $k$-mer tokenizers, this can be mitigated in $k$-mers with data augmentation using random offsets during pretraining (see Table 12).
5 Discussion
Our results demonstrate that pretraining masked language models on DNA barcode data, as exemplified by BarcodeBERT, is highly effective for arthropod species identification. Our pretrained model performs well on various downstream tasks common in biodiversity analyses, such as taxonomic classification, clustering, and similarity searches. BarcodeBERT excels in these tasks because it efficiently uses hardware acceleration, enabling it to scale effectively for large datasets, while being faster (55x) than alignment-based approaches in species level classification. By systematically evaluating tokenization and masking strategies, we also provide actionable insights for the pretraining of DNA-specific foundation models.
Despite its strengths, BarcodeBERT has some limitations. For instance, its training data may have taxonomic and geographical biases. In particular, the model is trained exclusively on invertebrate species from Canada, potentially limiting its applicability in global studies. In this context, the BOLD dataset, now comprising more than 16 million DNA barcodes from a wide geographical distribution (Ratnasingham & Hebert, 2007), represents a wealth of untapped data that could address these biases. Future work should incorporate more diverse datasets to develop robust, globally scalable self-supervised models for taxonomic classification. The methodology and findings we present should be broadly applicable to barcode regions for other kingdoms, such as the ITS region for fungi (Schoch et al., 2012), however the method is yet to be validated beyond the COI barcode region used for Animalia.
Lastly, while genomic sequences longer than barcodes could offer deeper insights for specialized phylogenetic analyses, the quadratic time complexity of transformer models limits their applicability to such sequences. Future work should include more parameter-efficient architectures such as Structured State Space Models, which scale sub-quadratically with sequence length (Gu et al., 2022). These architectures may be pretrained on specific barcode datasets to provide a more efficient alternative for applications involving longer sequences.
6 Conclusions
BarcodeBERT leverages 1 million DNA barcodes with partial taxonomic annotations to outperform state-of-the-art foundation models in genus-level and species-level classification tasks. Notably, BarcodeBERT matches the high accuracy of the alignment-based classification tool BLAST in species classification, while being 55 times faster and more scalable. In addition, our extensive analysis of pretraining strategies provides actionable insights for building customized DNA language models for large-scale taxonomic classification.
Overall, BarcodeBERT’s performance demonstrates how transformer-based architectures can be successfully customized to overcome the challenges of genomic biodiversity data, for effective DNA barcode identification and classification. Lastly, not being limited to a specific dataset or barcode region, our model is highly amenable to future applications, to global datasets or barcode regions in other kingdoms of life.
7 Competing interests
No competing interest is declared.
8 Author contributions statement
PMA curated the data, and partitioned it with assistance from GWT, DS, MS. PMA, SCL, MS, and NS implemented BarcodeBERT. PMA ran the DNA baseline experiments. MS conducted the masking and loss penalty ablation studies. NS conducted tokenization ablation studies. ZMG and ATW conducted the multimodal retrieval learning experiments with PMA’s assistance. PMA, MS, NS, LK, and IZ authored the manuscript text and figures. SCL, GWT, AXC, DS, LK, and JBH provided guidance on experimental design and edited the manuscript. All authors reviewed the manuscript.
9 Acknowledgments
We acknowledge the support of the Government of Canada’s New Frontiers in Research Fund [NFRFT-2020-00073]. This research was supported, in part, by the Province of Ontario and the Government of Canada through the Canadian Institute for Advanced Research (CIFAR), and companies sponsoring [^2] https://vectorinstitute.ai/partnerships/current-partners/ the Vector Institute. GWT is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada Research Chairs program, and the Canada CIFAR AI Chairs program. LK is supported by NSERC Discovery Grant RGPIN-2023-03663. DS is supported by the Canada First Research Excellence Fund through the University of Guelph’s “Food From Thought” program [Project 000054].
Appendix A Dataset Description
After pre-processing (see main text), the final dataset contained 965,289 samples. The dataset was divided into three main partitions: Pretrain, Seen, and Unseen. The Seen partition was subdivided into supervised training, validation, and testing subsets. Each partition serves distinct experimental purposes, with the Unseen partition designed to mimic real-world scenarios where models encounter sequences from previously unobserved species.
Table 6 summarizes the composition of each partition, detailing the number of unique records across taxonomic categories, from phylum to species. It also includes the Barcode Index Number (BIN), a species-level proxy formed by clustering similar DNA barcode sequences into a molecular taxonomic unit. The Pretrain partition contains 893,744 sequences from 15 phyla, while the Seen and Unseen partitions are more specific, containing sequences primarily from phylum Arthropoda. The Unseen partition contains 4,278 barcodes from 1,826 species absent from the training and validation subsets.
Table 6: Number of unique records across taxonomic categories in each partition.
Partition Phylum Class Order Family Genus Species DNA barcode BIN
Pretrain 15 49 176 1188 6679 14794 893744 62489
Test 1 1 13 161 500 1653 13460 2125
Train 1 1 13 161 500 1653 47086 2430
Validation 1 1 13 161 500 1653 6721 1954
Unseen 1 1 13 161 500 1826 4278 1885
To verify the overlap between partitions, Tables 7 and 8 provide pairwise comparisons of shared species and genera, respectively. As expected, there is no overlap in species between the Unseen partition and other subsets, ensuring the appropriateness of the Unseen partition for evaluating generalization. That being said, some genera are shared across partitions, reflecting the hierarchical taxonomic structure of the data. All the subsets in the Seen partition share all 500 genera that are also present in the Unseen partition. The Pretrain partition shares 467 of these genera with each of the others.
Table 7: Pairwise comparison of species overlap between partitions, with each cell representing the number of species shared between two partitions. The total number of unique species within a partition is given along the diagonal.
Pretrain Unseen Train Test Validation
Pretrain 14794 0 809 809 809
Unseen 0 1826 0 0 0
Train 809 0 1653 1653 1653
Test 809 0 1653 1653 1653
Validation 809 0 1653 1653 1653
Table 8: Pairwise comparison of genus overlap between partitions, with each cell representing the number of genera shared between two partitions. The total number of unique genera within a partition is given along the diagonal.
Pretrain Unseen Train Test Validation
Pretrain 6679 467 467 467 467
Unseen 467 500 500 500 500
Train 467 500 500 500 500
Test 467 500 500 500 500
Validation 467 500 500 500 500
Table 9 details the number of unique records across taxonomic ranks for each phylum in the Pretrain partition. Arthropoda dominates this partition, accounting for over 95% of the sequences, followed by smaller contributions from other phyla like Mollusca and Annelida. Table 10 highlights the percentage of labelled records for each phylum, showing significant gaps in taxonomic annotations at the species and genus levels, particularly for Arthropoda, where less than 40% of the sequences are labelled at the species level.
Table 9: The distribution of barcode sequences across taxonomic ranks for each phylum used in the Pretrain partition.
Phylum Class Order Family Genus Species BIN
Annelida 2 16 48 150 329 516
Arthropoda 14 67 929 6211 13991 61328
Brachiopoda 1 2 2 2 2 2
Bryozoa 3 3 3 2 2 4
Chordata 5 18 37 67 89 102
Cnidaria 4 10 24 25 24 46
Echinodermata 5 17 26 43 74 79
Hemichordata 1 1 1 2 1 2
Mollusca 6 30 97 162 271 372
Nematoda 2 5 10 5 2 8
Nemertea 3 2 5 5 5 22
Platyhelminthes 0 0 0 0 0 1
Porifera 1 3 4 4 3 5
Priapulida 1 1 1 1 1 1
Tardigrada 1 1 1 0 0 1
Table 10: Percentage of labelled records across taxonomic ranks for each phylum in the Pretrain partition.
Phylum Class Order Family Genus Species BIN
Annelida 100.00 96.43 99.81 87.44 80.35 97.95
Arthropoda 100.00 100.00 99.97 66.91 35.05 99.37
Brachiopoda 100.00 100.00 100.00 100.00 90.00 90.00
Bryozoa 80.00 80.00 80.00 40.00 40.00 100.00
Chordata 100.00 99.65 98.62 98.27 94.81 96.89
Cnidaria 100.00 98.21 95.54 84.82 54.46 99.11
Echinodermata 100.00 100.00 100.00 100.00 100.00 100.00
Hemichordata 100.00 100.00 100.00 100.00 25.00 100.00
Mollusca 99.95 89.02 99.69 98.43 90.22 96.71
Nematoda 100.00 100.00 91.67 54.17 16.67 37.50
Nemertea 96.43 94.64 87.50 64.29 48.21 100.00
Platyhelminthes 0.00 0.00 0.00 0.00 0.00 100.00
Porifera 100.00 100.00 100.00 85.71 42.86 85.71
Priapulida 100.00 100.00 100.00 100.00 100.00 100.00
Tardigrada 100.00 100.00 100.00 0.00 0.00 100.00
Figure 4 illustrates the species frequency across partitions, sorted in decreasing order. The subsets in the Seen partition exhibit a relatively uniform distribution, while the Unseen partition contains a higher proportion of rare species. Figures 5 and 6 present sunburst plots for the supervised test subset and the Unseen partition, visualizing the taxonomic hierarchy from order to genus. These figures highlight the differences in taxonomic composition between partitions, with the supervised test subset showing a more balanced distribution compared to the Unseen partition, which emphasizes underrepresented genera and families.
(a) Train subset (Seen)
(b) Test subset (Seen)
(c) Validation subset (Seen)
(d) Unseen partition
(e) Pretrain partition
Figure 4: Frequency of species records across each partition, sorted in decreasing order.
Figure 5: Sunburst plot of taxonomic distribution in the test subset of the Seen partition, with each radial level representing a taxonomic rank from order to genus. An interactive version of this plot is available at Test_Seen.html
Figure 6: Sunburst plot of taxonomic distribution in the Unseen partition, with each radial level representing a taxonomic rank from order to genus. An interactive version of this plot is available at Unseen.html
The partitions provide a structured dataset that supports the evaluation of models in both closed- and open-world settings. It enables a robust assessment of models’ ability to classify known taxa and generalize to new species.
Appendix B Tokenization Strategies — Extra results
B.1 Variation in Length of Tokenized Sequences
Compared to the $k$-mer tokenizer, which generates a tokenized sequence proportional to the DNA sequence length, the length of the tokenized sequence produced by the BPE tokenizer is not directly determined by the nucleotide sequence length and can vary depending on the composition of the input sequence and the vocabulary size of the BPE tokenizer.
Table 11 shows the minimum, mean, and maximum lengths of tokenized sequences in DNA barcode pretraining data. Generally, BPE tokenizers with smaller vocabulary sizes tend to produce longer tokenized sequences, with the only exception being the DNABERT-2 BPE. According to Table 11, although the tokenizer trained on DNABERT-2 uses a vocabulary size of 4096, it generates relatively long tokenized sequences. This is because the dataset used to train DNABERT-2 differs from the DNA barcode dataset used to create other tokenizers.
Table 11: Comparison of tokenized sequence lengths in the DNA barcode pretraining data across different BPE tokenizers with varying vocabulary sizes.
Tokenizer vocabulary sizes ($v$) Min Mean Max
DNABERT-2 4096 48 126.95 398
BarcodeBERT 4096 20 77.17 388
1024 30 104.61 407
256 54 146.59 429
128 65 175.44 451
B.2 Sensitivity of Tokenizers to DNA Sequence Variations
There are two main ways to measure similarity between sequences. One is through the Hamming distance, which counts the number of mismatches between two strings, i.e., the number of positions where the strings do not share the same nucleotide. The other is the edit distance, also called Levenshtein distance, which is defined as the minimum number of string operations (substitutions, insertions, deletions) required to transform one string into the other. In comparative genomics, the edit distance is widely used to estimate evolutionary relationships, as it corresponds to a better model of true sequence similarity. However, the Hamming distance can be calculated in linear time, even in token space with larger vocabularies. Figure 7 illustrates how the Hamming distance varies with a single insertion operation versus a single substitution operation in both sequence space and token space for a pair of sequences. We include the two tokenization strategies considered in our study: $k$-mer and DNABERT-2 BPE tokenization. In both scenarios, the edit distance is equal to one, but just the BPE tokenizer can model this behaviour using the Hamming distance as a proxy.
(a) Substitution
(b) Insertion
Figure 7: Illustration of the differences between the Hamming distances in sequence and token space using $k$-mer and DNABERT-2 BPE tokenization strategies for two different string operations where the edit distance between a pair of sequences is equal to one.
To better understand the relationships between string distances and their practical utility, we considered 1000 different genera in the Pretraining partition and sampled a pair of sequences from two different species within each genus. Figure 8 illustrates how the edit distance varies with the Hamming distance in the experiment. Two salient regions are evident in the plot. In the “aligned region”, the edit distance closely follows the Hamming distance as the number of mismatches is not due to alignment issues. In contrast, the “non-aligned region” potentially involves scenarios where sequences have been shifted, rearranged, or otherwise misaligned, leading to increased apparent mismatches when aligned positionally. These mismatches can be resolved with fewer string operations, resulting in a considerably lower edit distance compared to the Hamming distance.
Figure 8: Relationship between edit distance and Hamming distance for pairs of sequences sampled from different species across 1000 genera. The plot highlights two regions: the “aligned region”, where the edit distance and Hamming distance are closely matched, and the “non-aligned” region, where misalignments result in greater Hamming distances relative to the edit distance.
The Hamming distance in the token space of an optimal tokenizer should closely follow the behaviour of the edit distance in both regions. This is an indicator of how much the tokenizer is helping in modelling the true sequence similarity. To study the behaviour of both tokenization strategies, we repeated the experiment with 1000 random genera from the pretraining dataset and visualized a scatter plot of the Hamming distance in token space versus the Hamming distance in sequence space for these sequences for both BPE and $k$-mer tokenizers.
Figure 9: Comparison of Hamming distances in token space versus sequence space for pairs of sequences sampled from different species across 1000 genera. For both tokenizers, each dot in the plot represents the distances between sequences in a specific genus.
As illustrated in Figure 9, in the “aligned region”, the $k$-mer tokenizer closely corresponds to the Hamming distance scaled by a factor of $1/k$. However, the BPE tokenizer is sensitive to minor changes in the sequences, possibly due to substitutions that lead to unknown tokens, causing the tokenizer to split these unknown tokens into single nucleotides. Even when the Hamming distance between two DNA sequences is small, their tokenized representations can differ significantly with BPE tokenization compared to $k$-mer tokenization.
On the other hand, the BPE tokenizer is more robust to frame shifts, following the behaviour of the edit distance in the “non-aligned region” of the plot. The $k$-mer tokenizer struggles in this region, as large Hamming distances can sometimes result from small frame shifts in DNA sequences, which is also illustrated in Figure 7. This limitation can be addressed by incorporating data augmentation during the pretraining phase to increase the model’s robustness to frame shifts by randomly offsetting the input sequence.
Table 12 shows the effect of pretraining with a random offset augmentation on the downstream genus-level accuracy for unseen species when using $k$-mer versus BPE tokenization with the BarcodeBERT model (4 layers, 4 attention heads). The $k$-mer length is set to $k=4$, and for both tokenizers the offset is randomly selected between 0 and 3 (inclusive). We use a weight of 1 for the substitution component of the loss function ($w_{s}=1$), a substitution token ratio of 50% ($r_{s}=50\%$), and the proportion of substitution tokens assigned to [MASK] set to 1 ($r_{\texttt{[MASK]}}=1$). As shown in Table 12, random offset augmentation improves the genus-level accuracy of the $k$-mer-based model by 4.27 percentage points, but marginally decreases the performance of the model trained with the BPE tokenizer.
Table 12: Effect of augmenting with random offsets to the DNA sequences before tokenization with either $k$-mer or BPE tokenizers. In all experiments, we used a BarcodeBERT model with 4 layers and 4 attention heads, fixed weight for the substitution component of the loss function ($w_{s}\!=\!1$), substitution token ratio ($r_{s}\!=\!50\%$) and substitution token proportion assigned to [MASK], ($r_{\texttt{[MASK]}}\!=\!1$).
Tokenizer Random Offset Genus-level acc (%) of unseen species with $1$-NN probe
k-mer 74.20 ${\mathord{\uparrow}}4.27$
78.47 $\mathord{\uparrow}4.27$
BPE 69.85 ${\mathord{\downarrow}}1.29$
68.56 $\mathord{\downarrow}1.29$
Appendix C Masking Strategy
In this extra experiment, we adapted our masking strategy to mimic BERT’s original methodology (Devlin et al., 2019). BERT addresses the difference in token distribution between pretraining and fine-tuning by employing a masking strategy whereby 80% of substitution tokens are replaced by [MASK] tokens, 10% are replaced with random tokens, and 10% remain unchanged. This approach ensures more robust embeddings during testing, where masked tokens are absent (Devlin et al., 2019).
To incorporate this methodology into BarcodeBERT, we define the following terms: $r_{\texttt{[MASK]}}$ as the proportion of substitution tokens assigned to the [MASK] token and $r_{\texttt{[RAND]}}=1-r_{\texttt{[MASK]}}$ as the proportion of the substitution tokens assigned a random valid token (all tokens except the special tokens). We explored various ratios for token replacement and three different values for the substitution loss penalty $w_{s}$. In the first case, based on the results of our experiments, where $w_{s}=1$ had the best performance, we kept $w_{s}=1$ and adjusted $r_{\texttt{[MASK]}}$. In the second case, we set $w_{s}$ to 0.95 and in the third case, $w_{s}$ was set to 0.90, to closely replicate BERT’s original strategy that keeps 10% of the tokens unchanged. Note that in all experiments $r_{\texttt{[RAND]}}$ was set to $1-r_{\texttt{[MASK]}}$. In this study, we used the best configuration of 4 layers and 4 attention heads, $k=4$, and $r_{s}=50\%$. Figure 10 presents the accuracy of these experiments for genus-level 1-NN probing on unseen species. The results show that for $w_{s}=1.0$, the best accuracy is 78.47% with $r_{\texttt{[MASK]}}=1.0$, for $w_{s}=0.95$ the best accuracy is 76.85% with $r_{\texttt{[MASK]}}=0.9$, and for $w_{s}=0.9$, $r_{\texttt{[MASK]}}=0.5$ gives the best accuracy of 78.14%, which improves the accuracy by 1% compared to the case where $w_{s}=0.9$ and $r_{\texttt{[MASK]}}=1.0$. Our results demonstrate that adopting BERT’s masking strategy did not enhance the performance of BarcodeBERT, indicating that maintaining $r_{\texttt{[MASK]}}=1$ is the optimal configuration.
Figure 10: Genus-level accuracy for 1-NN probing of unseen species across different values of $r_{\texttt{[MASK]}}$, $r_{\texttt{[RNAD]}}$, and $w_{s}$. Experiments were conducted using the optimal configuration: 4 layers, 4 attention heads, $k$= 4, with substitution token ratio ($r_{s}=50\%$).
Appendix D Methodology — Extra details
To assess BarcodeBERT’s performance, we employed a combination of alignment-based methods, transformer-based models, and clustering algorithms. Here, we describe the parameters and hardware used in each methodology for reproducibility.
D.1 BLAST
We used BLAST (Altschul et al., 1990) for species-level classification to align query sequences from the test subset of the Seen partition against a nucleotide database built from the training subset of the Seen partition. To ensure high-quality alignments, we restricted the output to the top-scoring hit per query and applied strict thresholds of at least 80% query coverage and 88% sequence identity. For genus-level classification, we used BLAST to align query sequences from Unseen partition against the same database built from the training subset of the Seen partition. We did not enforce a minimum coverage or sequence identity in this experiment as the expected sequence similarity at the genus level was lower than at the species level. We ran the experiments on an Intel(R) Xeon(R) CPU @ 2.20GHz using 4 CPU threads to maximize computational performance. The full commands used to run the experiments are as follows:
    blastn -query supervised_test.fas -db train.fas
            -out results_supervised_top.tsv
            -max_target_seqs 1 -qcov_hsp_perc 80
            -perc_identity 88 -num_threads 4
            -outfmt 6
    blastn -query unseen.fas -db train.fas
            -out results_unseen.tsv -max_target_seqs 1
            -num_threads 4 -outfmt 6
D.2 Baseline Models
For evaluation, we utilized the respective Pretrained models from Huggingface ModelHub, specifically:
D.3 Pretraining
BarcodeBERT was pretrained for 35 epochs using the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $2\times 10^{-4}$, a batch size of 128, and a OneCycle learning rate scheduler (Smith & Topin, 2017). The pretraining process utilized four NVIDIA V100 GPUs and required approximately 36 hours to complete for each experiment executed. To examine the impact of pretraining, we also trained a model from scratch on the training subset of the Seen partition without any pretraining.
D.4 Fine-tuning
All baseline models and BarcodeBERT were fine-tuned for 35 epochs on the supervised training subset of the Seen partition. For larger models, a batch size of 32 was used, while smaller models (CNN, HyenaDNA and BarcodeBERT) were trained with a batch size of 128. The AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $1\times 10^{-4}$ was employed, coupled with the OneCycle learning rate scheduler (Smith & Topin, 2017). We used a single NVIDIA V100 GPU for the fine-tuning process, completing within 18 hours for most models. Notably, HyenaDNA required significantly less time to fine-tune compared to other models, as a consequence of its lightweight architecture.
D.5 Linear probe training
A linear classifier is applied to the embeddings generated by all the pretrained models for species-level classification. The models’ parameters are learned using stochastic gradient descent with a constant learning rate of 0.01, momentum $\mu=0.9$, and weight $\lambda=1\times 10^{-5}$.
D.6 Zero-shot clustering
We evaluated the models’ ability to group sequences without supervision using the framework from Lowe et al. (2024). Embeddings were extracted from the pretrained encoders and reduced to 50 dimensions using UMAP (McInnes et al., 2018) to enhance computational efficiency while preserving data structure. These reduced embeddings were clustered with Agglomerative Clustering (L2 distance, Ward’s linkage), using the number of true species as the target number of clusters. Clustering performance was assessed with adjusted mutual information (AMI) to measure alignment with ground-truth labels.